OpenStudy (amanda0618):

write an equation in point slope form for the line through the given point with the given slope (7,-2);m=5

OpenStudy (kamibug):

point slope form is y−y1=m(x−x1) so plug in the slope and point 7 is x_1 and -2 is y_1

OpenStudy (kamibug):

\(\huge\color{navy}{Welcome~to~Openstudy!~:)}\)

OpenStudy (amanda0618):

could you please type in point slope form again? it vanished from my screen @KamiBug

OpenStudy (anonymous):

(y-y1)=m(x-x1)